OpenStudy (anonymous):

a certain medicine is given in an amount proportional to a patients body weight.suppose a patient weighed 100 pounds requires 125 milligrams of medicine,what is the weight of a patient who requires 300 milligrams of medicine?

OpenStudy (kira_yamato):

Let weight = W pounds, amount of medicine = m mg W is directly proportional to m W = km, where k is constant of proportionality, k > 0 When W = 100, m = 125 100 = k(125) k = 0.8 W = 0.8m When m = 300, W = 0.8(300) = 240 Weight = 240 pounds

OpenStudy (anonymous):

240 pounds

OpenStudy (anonymous):

KIRA!!!

OpenStudy (anonymous):

here is a simpler way...300/125 = 2.4 x 100 = 240 (: Hope for a medal :3